whitemonsterbunny17 (whitemonsterbunny17):

More help drawing special polygons please? Thanks!! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

Sketch the polygon described. If no such polygon exists, write "not possible". A regular pentagon, one of whose angles has measure 120.

whitemonsterbunny17 (whitemonsterbunny17):

This might be a reeeeeally dumb question, but what is a pentagon?? D:

OpenStudy (sleepyjess):

not possible a regular pentagon has angle measures of 108 pentagon = 5 sides

ganeshie8 (ganeshie8):

a pentagon with one angle measuring 120 is not a regular

OpenStudy (anonymous):

A special polygon is a polygon that is special to one's soul. So just draw a polygon that your heart desires :DDD

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

also have u found answer for ur previous question about equilateral quadrilateral ?

whitemonsterbunny17 (whitemonsterbunny17):

Ohhh, okay ^.^ Btw, how do you know all the total measures of different shapes? And yes, I did get the answer @ganeshie8

OpenStudy (sleepyjess):

if you just google regular pentagon angles it tells you lol

ganeshie8 (ganeshie8):

yes google tells everything :)

ganeshie8 (ganeshie8):

i usually work it with external angle

ganeshie8 (ganeshie8):

all u need to remember is this : 1) sum of external angles of ANY polygon = 360 2) internal angle + external angle = 180

ganeshie8 (ganeshie8):

for a pentagon : \[\large x+x+x+x+x = 360 \]

ganeshie8 (ganeshie8):

solving that gives you \[x = 72\]

OpenStudy (sleepyjess):

a pentagon adds up to 540 @ganeshie8

ganeshie8 (ganeshie8):

btw, \(x\) is external angle

OpenStudy (sleepyjess):

oh ok nvm

ganeshie8 (ganeshie8):

subtract it from 180 to get the internal angle : \[\large \text{internal angle} = 180-72 = 108\]

whitemonsterbunny17 (whitemonsterbunny17):

Ooooohhh, okay that makes sense!! I knew that the exterior measure should be 360, but I didn't know that goes for all polygons xD

ganeshie8 (ganeshie8):

that goes for ALL polygons because all polygons are closed (they complete a circle )

ganeshie8 (ganeshie8):